Substantive Test

Assume that you are now performing substantive tests on receivables and you want to perform account receivable confirmations. The client’s recorded book value is $2,300,000 made up of 2,650 customers. You have established a tolerable misstatement of $75,000 for accounts receivable. You have selected 200 accounts worth a total of $235,000 to confirm and after sending out the confirmations, four are returned showing the following overstatements.

Book Value

Audit value

Sonic Trading            

12,800

10,350

Bazaar Enterprises      

8,300

6,500

Millers Co                    

18,800

14,150

Mendez Bros.               

4,925

3,200

2a.  What is your sample misstatement?

2b.  What is  your projected misstatement using the ratio estimation method?

2.c What do you conclude about the fairness of the accounts receivable account?

2d. Using the difference estimation method, calculate your projected misstatement

2e. What do you conclude about the fairness of the accounts receivable account?
Answer #1
Book Value Audit value Sample misstatement(Book Value-Audit Value)
Sonic Trading             12,800 10,350 2,450
Bazaar Enterprises       8,300 6,500 1,800
Millers Co                     18,800 14,150 4,650
Mendez Bros.                4,925 3,200 1,725
Total 44,825 34,200 10,625
2a.Net Sample misstatement= $ 10625
2.b.Projected misstatement using the ratio estimation method
(Net sample misstatement/Book value of Sample)*Population book value
10625/44825*2300000=
545176
(overstatement)
ie. Misstatement ratio=10625/44825=23.70%
23.70% applied on total a/cs rec.=2300000*23.70%= approx. $ 545100
So,estimated audit value of accounts receivables,( eliminating overstatement error)will be
2300000-545176=
1754824
2.c. As this $   545176 is much > the tolerable amt. of $ 75000 --Accounts receivables are likely to be materially misstated.
2d.The projected misstatement using difference estimation method:
Projected misstatement=(Sample net misstatement/No.of Sample items)*No.of population items
ie. 10625/200*2650=
140781
(overstatement)
So,estimated audit value of accounts receivables, will be
2300000-140781=
2159219
2e.Accounts receivables are likely to be materially misstated.